Valley Bank has an average interest rate of 10% on %500 million of loans and securities, an average interest rate of 5% on $400 million of deposits and other borrowed funds, noninterest revenue of $12 million, noninterest expense of $17 million, and a provision for loan losses of $10 million. total shareholders equity is $50 million, total assets are $500 million and the tax rate is 30%. Prepare Valley's income statement in order to answer:

A. the net interest income

B. net income

C. return on asset
Answer #1

a.

Net Interest Income = 0.10*500 - 0.05*400 = 30 million

b.

Net Income = (30 + 12 - 17 - 10)(1 - 0.30)

Net Income = 10.5 million

c.

ROA = 10.5/500

ROA = 2.1%